Sam can cultivate a piece of land in 7hours while Ian can do the same work in 5hours. Find time they would take to cultivate the same piece of land when working together
Answer by Boreal(15235)

in 1 hour, Sam can do 1/7 and Ian can do 1/5
in x hours, (x/7)+(x/5)=1
multiply by 35 both sides
5x+7x=35
x=35/12
or 2 11/12 hours working together or 2h55m